# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v3

@@ Abstract.
This article is inspired from the work of M Krithika and P Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8]. We establish the existence of polynomials for given degree and cluster size over number fields which generalises a result of Perlis. We state the Strong cluster magnification problem and establish an equivalent criterion for that. We also discuss the notion of weak cluster magnification and prove some properties. We provide an important example answering a question about Cluster Towers. We introduce the concept of Root capacity and prove some of its properties. We also introduce the concept of unique descending and ascending chains for extensions and establish some properties and explicitly compute some interesting examples. Finally we establish results about all these phenomena under a particular type of base change. The article concludes with results about strong cluster magnification, unique chains and some properties of the ascending index for a field extension.
@@ 1. Introduction
In Sec. 2, we set up some notations and review the results by Perlis in [7], and Krithika-Vanchinathan in [8], [5].
In Sec. 3, we generalise a result of Perlis for number fields that also improves on the generalisation proved previously by Krithika and Vanchinathan in [5]. This generalisation, Thm. 3.1.1 establishes existence of polynomials for given degree and cluster size over number fields. In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very useful in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
@@
An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
@@
In Sec. 6, we introduce the concept of Root capacity. We begin the section with some observations about the group of automorphisms of finite extensions in Prop. 6.1.1 and Prop. 6.1.3 and its corollaries. Then we prove some properties of root capacity and conclude with Thm. 6.2.13.
@@
In Sec. 8, we establish results Thm. 8.1.1, 8.1.9, 8.2.2 and 8.3.1 about strong/weak cluster magnification, root capacity and unique chains, respectively under a particular type of base change. We prove results Thm. 8.4.1 and Thm. 8.4.3 about strong cluster magnification and unique chains.
The article concludes with Sec. 9 with some properties of the ascending index $t_{K}(L)$ defined in Thm. 7.2.1 in the context of unique ascending chain for an extension $L/K$. The ascending index has many properties similar to the cluster size $r_{K}(L)$ but has no immediate description in terms of roots of the minimal polynomial of $\alpha$ over $K$ when $L=K(\alpha)$. In Sec. 9, Prop. 9.0.3 establishes a base change property for $t_{K}(L)$, Prop. 9.0.4 establishes an analogue of Cluster Magnification Theorem (Thm. 2.2.3) for $t_{K}(L)$ and Thm. 9.0.5 is an analogue of Thm. 3.1.1 and establishes existence of an extension $L/K$ with prescribed degree and $t_{K}(L)$.
@@ 2.1. Root clusters
Let $s_{K}(f)$ be the number of distinct fields of the form $K(\alpha_{j})$, with $\alpha_{j}$ a root of $f$ in $K_{f}$ for all $1\leq j\leq{\rm deg}(f)$, i.e., $s_{K}(f)$ is the number of clusters of roots of $f$ in $K_{f}$.
The following result is proved in [7] and [8].
@@ 2.2. Cluster Magnification Theorem
We note the following basic fact from Galois theory.
@@
The following lemma can be deduced from [4, Lem. 1, Chap. 8.15] in combination with Remark 2.2.1.
@@
[Krithika, Vanchinathan] Let $K,f$ and $\alpha$ be as above. Let ${\rm deg}(f)=n>2$ over $K$ with cluster size $r_{K}(f)=r$. Assume that there is a Galois extension $F$ of $K$, say of degree $d$, which is linearly disjoint with $K_{f}$ over $K$. Then there exists an irreducible polynomial $g$ over $K$ of degree $nd$ with cluster size $rd$. ($F$ can be chosen to be $K(\beta)$ for some $\beta$ in $\bar{K}$ so that $K(\alpha,\beta)=K(\alpha+\beta)$ and the irreducible polynomial of $\alpha+\beta$ over $K$, has degree $nd$ with cluster size $rd$; $d$ is the magnification factor).
Reformulation : Let $K$ be a perfect field as above and $L/K$ be an extension of degree $n>2$ contained in $\bar{K}$ with cluster size $r_{K}(L)=r$. Let $F/K$ be any finite Galois extension of degree $d$ contained in $\bar{K}$, which is linearly disjoint with $\tilde{L}$ over $K$. Then the compositum $LF/K$ has degree $nd$ with cluster size $r_{K}(LF)=rd$. ($d$ is the magnification factor).
@@ 3.1. Existence of polynomials for given degree and cluster size over number fields
Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
@@
Let $K$ be a number field. We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([10, Def. 1.14 and Example 1.17]). In particular, ${\mathfrak{S}}_{n}$ occurs regularly over $K$. The result [10, Corollary 1.15] says that if a group occurs regularly over a hilbertian field, then it is realisable as a Galois group over that field. Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$.
The following lemma is the final proposition in Perlis [7]. We write the proof given by Perlis for the sake of completion.
@@
As a corollary, we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$. This corollary also holds for $G^{n}$ in place of $G$ for any $n\in\mathbb{N}$ as well as for $K^{\prime}$.
@@
Since $K$ is perfect and $K^{\prime}/K$ is finite, we have that $K^{\prime}/K$ is separable. Hence, $K^{\prime}/K$ has finitely many intermediate fields. Let $n$ be the number of these intermediate fields (including $K^{\prime}$ and $K$). Now $G^{n}$ is realisable over $K$ by assumption, say for $E/K$ Galois, we have ${\rm Gal}(E/K)\cong G^{n}$. We have normal subgroups $N_{i}=G\times G\times\dots\times{1}\times\dots\times G$ of $G^{n}$ for $1\leq i\leq n$ where the $i$th coordinate is trivial and there is no restriction in other coordinates. So $N_{i}\cong G^{n-1}$. Let $E_{i}$ be the subfield of $E$ corresponding to $N_{i}$, so $E_{i}/K$ is Galois with ${\rm Gal}(E_{i}/K)\cong G^{n}/N_{i}\cong G$.
Now for $i\neq j$, $E_{i}\cap E_{j}$ corresponds to subgroup generated by $N_{i}$ and $N_{j}$ which is $G^{n}$. Hence $E_{i}\cap E_{j}=K$. Suppose for some $i\neq j$ we have $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}$. Since $E_{i}\cap E_{j}=K$, we get $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$. Now suppose that all $E_{i}\cap K^{\prime}$ are distinct. Since we have $n$ intermediate fields of $K^{\prime}/K$, $E_{i}\cap K^{\prime}=K$ for some $i$. In either case we get an $i$ such that $E_{i}\cap K^{\prime}=K$. Hence ${\rm Gal}(E_{i}K^{\prime}/K^{\prime})\cong{\rm Gal}(E_{i}/K)\cong G$. This realises $G$ as a Galois group over $K^{\prime}$. By replacing $G$ with $G^{m}$ for any $m\in\mathbb{N}$ in the above argument, we can realise $G^{m}$ over $K^{\prime}$ for any $m$.
For the proof of the corollary, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$. We observe that $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$. Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$. ∎
@@
Since direct product of solvable groups is solvable, direct products $G^{i}$ for $i\in\mathbb{N}$ are solvable. By Shafarevich’s theorem ([9]), $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$. Hence by Lemma 3.1.4, $G$ is realisable as Galois group over number field $K$. By Lemma 3.1.3, there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$. This $f$ satisfies $r_{K}(f)=r$.
@@ 3.2. A simple lemma about $s_{K}(L)$
Let $K$ be a perfect field. Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$ (For a number field $K$, such $f$ always exists. See $r=1$ case in Proof of Thm. 3.1.1).
For $21\leq k\leq n-$ 2, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$. Let $g$ be the irreducible polynomial over $K$ for a primitive element of $L_{k}$. This polynomial has degree ${}^{n}P_{k}$ and cluster size $k!$.
@@ 4.1. Strong cluster magnification
Hence $K^{\prime}F/K^{\prime}$ is Galois and $L_{1}$ and $K^{\prime}F$ are linearly disjoint over $K^{\prime}$. Also $M=LF=LK^{\prime}F$ and hence we are done. The magnification factor is same since $[F:K]=[K^{\prime}F:K^{\prime}]$.
@@
    We have isomorphism $G^{\prime}\cong G\times R$. Hence $G=G_{0}R_{0}$ where $G_{0},R_{0}\subset G$ with $G_{0}\cong G\times 1$ and $R_{0}\cong 1\times R$ under above isomorphism. Furthermore, $\tilde{L}=\tilde{M}^{R_{0}}$ and $F=\tilde{M}^{G_{0}}$. Now if a subextension $L$ of $M$ is given, then $\tilde{L}$ is uniquely determined inside $\tilde{M}$. Thus $R_{0}$ is uniquely determined inside $G$ which implies that $G_{0}$ is uniquely determined inside $G$. Hence $F$ is uniquely determined from $L$ and $M$.
@@
An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
@@
Conversely, suppose $G^{\prime}\cong A\times B$ for nontrivial subgroups $A$ and $B$ and $H^{\prime}\cong A^{\prime}\times\{e\}$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$. We identify $G^{\prime}$ and $H^{\prime}$ with their images under the isomorphism. Now we check the three conditions of Def. 4.1.1 for $M/K$.
@@
The magnification is nontrivial since $B$ is nontrivial subgroup. ∎
@@
Let $M/K$ be Galois. Then $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(M/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ with $|A|>2$. If this happens then, $L/K$ is also Galois.
@@
We end this section with a result that the strong cluster magnification behaves well with respect to $K$-isomorphisms.
Proposition 4.1.9.
@@
Proof.
Suppose $M/K$ is obtained by strong cluster magnification from $L/K$. We have $L/K$, $n$, $r$, $F/K$, $d$ as above. Now we check the three conditions of Def. 4.1.1 for $M^{\prime}/K$.
  1. (1)
    $\sigma(L)\cong L$ has degree $n>2$ over $K$ with $r_{K}(\sigma(L))=r_{K}(L)=r$.
  2. (2)
    $F=\sigma(F)$ since $F/K$ is Galois. It is easy to prove that $\tilde{L}$ is Galois closure of $\sigma(L)$ in $\bar{K}$. $\tilde{L}$ and $F$ are linearly disjoint over $K$.
  3. (3)
    $\sigma(L)F=\sigma(M)=M^{\prime}$.
Alternatively, we can use the criterion in Thm. 4.1.6. The isomorphism between $M$ and $M^{\prime}$ extends to isomorphism between their Galois closures. Hence we also get an isomorphism between $G^{\prime},H^{\prime}$ and corresponding groups of $M^{\prime}$.

### v4

# Cluster Magnification, Root Capacity, Unique Chains, Base Change and Ascending Index

@@ Abstract.
This article is inspired from the work of M Krithika and P Vanchinathan in [5] and the work of Alexander Perlis in [7] and [8]. We establish the existence of polynomials for given degree and cluster size over number fields which generalises a result of Perlis. We state the Strong cluster magnification problem and establish an equivalent criterion for that. We also discuss the notion of weak cluster magnification and prove some properties. We provide an important example answering a question about Cluster Towers. We introduce the concept of Root capacity and prove some of its properties. We also introduce the concept of unique descending and ascending chains for extensions and establish some properties and explicitly compute some interesting examples. We establish results about all these phenomena under a particular type of base change and discuss some other related results about strong cluster magnification and unique chains. The article concludes with results about ascending index for a field extension which are analogous to results about cluster size.
@@
Keywords: Galois theory, Root clusters, Base change, Ascending Index
@@ 1. Introduction
In Sec. 2, we set up some notations and review the results by Perlis in [7], [8] and Krithika-Vanchinathan in [5].
In Sec. 3, we generalise a result of Perlis for number fields that also improves on the generalisation proved previously by Krithika and Vanchinathan in [5]. The generalisation Thm. 3.1.1 is as follows.
Theorem 1.0.1.
(Inverse Cluster Size Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
In the same section, we also present a simple lemma about number of clusters, Lemma 3.2.1, which is very useful in giving alternate proofs of results by Perlis and Krithika-Vanchinathan as well as in proving further results.
@@
An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial groups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
@@
In Sec. 6, we introduce the concept of Root capacity. We begin the section with some observations about the group of automorphisms of finite extensions in Prop. 6.1.1 and Prop. 6.1.3 and its corollaries. Then we prove some properties of root capacity in Propositions 6.2.2, 6.2.4, 6.2.6, 6.2.9 and 6.2.11. We conclude the section with Thm. 6.2.13 which is as follows. For notations, see Sec. 2 and Sec. 6.
Theorem 1.0.3.
Consider extensions $M/L/K$. If $M\cap\tilde{L}=L$ and $[M:L]=r_{K}(M)/r_{K}(L)$, then $M/L$ is Galois.
@@
In Sec. 8, we define notion of a particular type of base change and then establish results Thm. 8.1.2, 8.1.10, 8.2.2 and 8.3.1 about strong/weak cluster magnification, root capacity and unique chains, respectively under this base change. We prove results Thm. 8.4.1 and Thm. 8.4.3 about strong cluster magnification and unique chains.
The article concludes with Sec. 9 with some properties of the ascending index $t_{K}(L)$ defined in Thm. 7.2.1 in the context of unique ascending chain for an extension $L/K$. The ascending index has many properties similar to the cluster size $r_{K}(L)$ but has no immediate description in terms of roots of the minimal polynomial of $\alpha$ over $K$ when $L=K(\alpha)$. In Sec. 9, Prop. 9.0.3 establishes a base change property for $t_{K}(L)$. Thm. 9.0.4 establishes an analogue of Cluster Magnification Theorem (Thm. 2.2.3) for $t_{K}(L)$ which is as follows. For notations, see Sec. 7.
Theorem 1.0.4.
(Ascending Index Magnification Theorem) Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then
$t_{K}(M)=d\ t_{K}(L)\ \text{and}\ u_{K}(M)=u_{K}(L).$
Finally, Thm. 9.0.5 is an analogue of Thm. 3.1.1 and is as follows.
Theorem 1.0.5.
(Inverse Ascending Index Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
@@ 2.1. Root clusters
Let $s_{K}(f)$ be the number of distinct fields of the form $K(\alpha_{j})$, with $\alpha_{j}$ a root of $f$ in $K_{f}$ for all $1\leq j\leq{\rm deg}(f)$. The following result is proved in [7] and [8].
@@
By the proof of above theorem it follows that $s_{K}(f)$ is the number of clusters of roots of $f$ in $K_{f}$.
@@ 2.2. Cluster Magnification Theorem
We note the following basic fact from Galois theory.
@@
The following lemma can be deduced from [3, Lem. 1, Chap. 8.15] in combination with Remark 2.2.1.
@@
(Krithika, Vanchinathan) Let $K,f$ and $\alpha$ be as above. Let ${\rm deg}(f)=n>2$ over $K$ with cluster size $r_{K}(f)=r$. Assume that there is a Galois extension $F$ of $K$, say of degree $d$, which is linearly disjoint with $K_{f}$ over $K$. Then there exists an irreducible polynomial $g$ over $K$ of degree $nd$ with cluster size $rd$. ($F$ can be chosen to be $K(\beta)$ for some $\beta$ in $\bar{K}$ so that $K(\alpha,\beta)=K(\alpha+\beta)$ and the minimal polynomial of $\alpha+\beta$ over $K$, has degree $nd$ with cluster size $rd$; $d$ is the magnification factor).
Reformulation : Let $K$ be a perfect field as above and $L/K$ be an extension of degree $n>2$ contained in $\bar{K}$ with cluster size $r_{K}(L)=r$. Let $F/K$ be any finite Galois extension of degree $d$ contained in $\bar{K}$, which is linearly disjoint with $\tilde{L}$ over $K$. Then the compositum $LF/K$ has degree $nd$ with cluster size $r_{K}(LF)=rd$ ($d$ is the magnification factor).
@@ 3.1. Existence of polynomials for given degree and cluster size over number fields
(Inverse Cluster Size Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an irreducible polynomial over $K$ of degree $n$ with cluster size $r$.
@@
Let $K$ be a number field. We know that the group ${\mathfrak{S}}_{n}$ occurs regularly over every field ([10, Def. 1.14 and Example 1.17]). In particular, ${\mathfrak{S}}_{n}$ occurs regularly over $K$. The result [10, Corollary 1.15] says that if a group occurs regularly over a hilbertian field, then it is realisable as a Galois group over that field. Hence finally we conclude that ${\mathfrak{S}}_{n}$ is realisable as Galois group over $K$. ∎
The following lemma is the final proposition in Perlis [7]. We write the proof given by Perlis for the sake of completeness.
@@
Since $K$ is perfect and $K^{\prime}/K$ is finite, we have that $K^{\prime}/K$ is separable. Hence, $K^{\prime}/K$ has finitely many intermediate fields. Let $n$ be the number of these intermediate fields (including $K^{\prime}$ and $K$). Now $G^{n}$ is realisable over $K$ by assumption, say for $E/K$ Galois, we have ${\rm Gal}(E/K)\cong G^{n}$.
We have normal subgroups $N_{i}=G\times G\times\dots\times{1}\times\dots\times G$ of $G^{n}$ for $1\leq i\leq n$ where the $i$th coordinate is trivial and there is no restriction in other coordinates. So $N_{i}\cong G^{n-1}$. Let $E_{i}$ be the subfield of $E$ corresponding to $N_{i}$, so $E_{i}/K$ is Galois with ${\rm Gal}(E_{i}/K)\cong G^{n}/N_{i}\cong G$.
Now for $i\neq j$, $E_{i}\cap E_{j}$ corresponds to subgroup generated by $N_{i}$ and $N_{j}$ which is $G^{n}$. Hence $E_{i}\cap E_{j}=K$. Suppose for some $i\neq j$ we have $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}$. Since $E_{i}\cap E_{j}=K$, we get $E_{i}\cap K^{\prime}=E_{j}\cap K^{\prime}=K$. Now suppose that all $E_{i}\cap K^{\prime}$ are distinct. Since we have $n$ intermediate fields of $K^{\prime}/K$, $E_{i}\cap K^{\prime}=K$ for some $i$. In either case we get an $i$ such that $E_{i}\cap K^{\prime}=K$. Hence ${\rm Gal}(E_{i}K^{\prime}/K^{\prime})\cong{\rm Gal}(E_{i}/K)\cong G$. This realises $G$ as a Galois group over $K^{\prime}$. By replacing $G$ with $G^{m}$ for any $m\in\mathbb{N}$ in the above argument, we can realise $G^{m}$ over $K^{\prime}$ for any $m$. ∎
Remark 3.1.5.
In the above proof, we observe that for any $n$ we have some $E/K$ Galois with Galois group $G^{n}$ and $N_{i}$ normal subgroups of $G^{n}$ and $E_{i}$ subfield of $E$ corresponding to $N_{i}$. We observe that $N_{i}$ are not conjugate to each other in $G^{n}$ and they pairwise generate $G^{n}$. Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$ with $G$ as Galois group of each $E_{i}/K$. Thus we conclude the following.
Let $K$ be a perfect field and $K^{\prime}/K$ be a finite extension and suppose for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$. Then we get arbitrarily large finite families of Galois extensions of $K$ inside a fixed $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$. This statement also holds for $G^{m}$ in place of $G$ for any $m\in\mathbb{N}$ as well as for $K^{\prime}$ in place of $K$.
@@
Since direct product of solvable groups is solvable, direct products $G^{i}$ for $i\in\mathbb{N}$ are solvable. By Shafarevich’s theorem ([9]), $G^{i}$ for $i\in\mathbb{N}$ are realisable as Galois groups over $\mathbb{Q}$. Hence by Lemma 3.1.4, $G$ is realisable as Galois group over number field $K$. By Lemma 3.1.3, there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$. This $f$ satisfies $r_{K}(f)=r$. ∎
@@ 3.2. A simple lemma about $s_{K}(L)$
Let $K$ be a perfect field. Let $f$ over $K$ be irreducible of degree $n$ with Galois group ${\mathfrak{S}}_{n}$. For $21\leq k\leq n-$ 2, let $L_{k}$ be an extension of $K$ obtained by adjoining any $k$ roots of $f$ in $\bar{K}$. Let $g$ be the minimal polynomial over $K$ for a primitive element of $L_{k}$. This polynomial has degree ${}^{n}P_{k}$ and cluster size $k!$.
@@ 4.1. Strong cluster magnification
Hence $K^{\prime}F/K^{\prime}$ is Galois and $L_{1}$ and $K^{\prime}F$ are linearly disjoint over $K^{\prime}$. Also $M=LF=LK^{\prime}F$ and hence we are done. The magnification factor is same since $[F:K]=[K^{\prime}F:K^{\prime}]$. ∎
@@
    We have isomorphism $G^{\prime}\cong G\times R$. Hence $G^{\prime}=G_{0}R_{0}$ where $G_{0},R_{0}\subset G^{\prime}$ with $G_{0}\cong G\times 1$ and $R_{0}\cong 1\times R$ under above isomorphism. Furthermore, $\tilde{L}=\tilde{M}^{R_{0}}$ and $F=\tilde{M}^{G_{0}}$. Now if a subextension $L$ of $M$ is given, then $\tilde{L}$ is uniquely determined inside $\tilde{M}$. Thus $R_{0}$ is uniquely determined inside $G^{\prime}$ which implies that $G_{0}$ is uniquely determined inside $G^{\prime}$. Hence $F$ is uniquely determined from $L$ and $M$.
@@
An extension $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(\tilde{M}/K)\cong A\times B$ for nontrivial groups $A$ and $B$ and ${{\rm Gal}}(\tilde{M}/M)\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$.
@@
Conversely, suppose $G^{\prime}\cong A\times B$ for nontrivial groups $A$ and $B$ and $H^{\prime}\cong A^{\prime}\times\{e\}$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$. We identify $G^{\prime}$ and $H^{\prime}$ with their images under the isomorphism. Now we check the three conditions of Def. 4.1.1 for $M/K$.
@@
The magnification is nontrivial since $B$ is nontrivial group. ∎
@@
Let $M/K$ be Galois. Then $M/K$ is obtained by nontrivial strong cluster magnification from some subextension $L/K$ if and only if ${{\rm Gal}}(M/K)\cong A\times B$ for nontrivial groups $A$ and $B$ with $|A|>2$. If this happens then, $L/K$ is also Galois.
@@
The following says, strong cluster magnification behaves well with respect to $K$-isomorphisms.
Remark 4.1.9.